Circular No. 5928

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 S. Benetti, A. Clocchiatti, and J. C. Wheeler, University of Texas,

report: “A fully reduced spectrogram (range 400-920 nm, resolution 0.9 nm) of SN 1993J, obtained on Jan. 17.5 UT at the 2.1-m telescope of McDonald Observatory, shows three strong emission lines and a very blue continuum: [O I] at 630.0 and 636.4 nm (integrated flux 5.8 × 10^{−13} erg s^{−1} cm^{−2} with the same fine structure as reported on IAUC 5847), Ca II] at 729.1 and 732.4 nm (flux 2.9 × 10^{−13} erg s^{−1} cm^{−2}), and a remarkably strong Mg I] at 457.1 nm (flux 1.3 × 10^{−13} erg s^{−1} cm^{−2}). The lines have sharp blue edges and fine structure. The Ca II infrared triplet is weak (flux 1.1 × 10^{−13} erg s^{−1} cm^{−2}) in comparison with typical type-II supernovae at the same phase. [Fe II] at 715.5 nm (flux 0.4 × 10^{−13} erg s^{−1} cm^{−2}) and Fe II at 492.4, 501.8, and 516.9 nm are also present. Narrow, weak circumstellar Hα emission (FWHM = 1.4 nm, barely resolved) with maximum at 655.8 nm, already present in a McDonald Observatory spectrum of 1993 Nov. 12, is broader and more blue-shifted (−230 km/s) than the circumstellar Hα emission observed just after outburst. Wide Hα emission with a sharp red edge extends to the red from the base of the [O I] doublet at 630.0 and 636.4 nm. Gaussian deconvolution yields FWHM = 40.5 nm centered at 656.6 nm. This feature may represent a shell of hydrogen expanding at (half-width) velocity of 10 000 km/s.”

GRO J0422+32 C. Chevalier and S. A. Ilovaisky, Observatoire de Haute-Provence,

write: “Time-resolved CCD photometry of this object obtained on 1993 Dec. 17/18 and 21/22 with the 1.2-m Haute-Provence telescope shows a light curve consistent with the period and phase determined from 21 nights in 1993 Jan. and Feb. (IAUC 5692, 5734). Assuming a double-wave light curve, the period is 0.4243 day with the peak-to-peak amplitude near 0.05 mag in V. The average brightness on these two nights was V = 15.5 and 15.8, respectively.”

NOVA CASSIOPEIAE 1993 Further photoelectric photometry by H. Mikuž, Ljubljana, Slovenia

(cf. IAUC 5920): Jan. 15.77 UT, V = 7.49; 19.76, 7.73; 22.82, 7.67.

1994 January 28 Daniel W. E. Green